# Unraveling the Age Estimation Puzzle: Comparative Analysis of Deep Learning Approaches for Facial Age Estimation

### v1

@@ 3.2 Model Architecture
Multiple different backbone architectures can be found in the age estimation literature. Among these architectures, VGG16 [13, 27, 26, 22, 17, 9] and ResNet-50 [3, 2, 19] stand out as the most common choice. We have chosen to use the ResNet-50 [14] as baseline backbone architecture and experiment with different architectures in Sec. 4.3.
3.3 Method Details
Note that for all methods which model the posterior distribution $p(y|x)$, namely (i) cross-entropy (ii) DLDL [12] (iii) DLDL-v2 [13] (iv) SORD [9] (v) Mean-Variance loss [22] (vi) Unimodal loss [17] , we use optimal plugin Bayes predictor for MAE loss, i.e., we predict $\arg\min_{y}\mathbb{E}_{\hat{y}\sim f(x)}[|y-\hat{y}|]$. For regression, we use the absolute error as the loss function.
Training Details
We utilize the Adam optimizer with the parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$. For pre-training on the IMDB-WIKI dataset, we set the learning rate to $\alpha=10^{-3}$ and train the model for a total of 100 epochs. For fine-tuning on the remaining datasets we reduce the learning rate to $\alpha=10^{-4}$ and train the model for 50 epochs. We use a batch size of 100. The best model is selected based on the MAE metric computed on the validation set. We utilize two data augmentations during training, (i) horizontal mirroring (ii) cropping out an 80% to 100% portion of the bounding box and resizing it to the model input shape.
3.4 Data
The main challenge in comparing different age estimation methods lies in the fact that published results often utilize different data splits. Even when using a consistent data-splitting protocol, such as random or subject-exclusive, the results can differ significantly (i.e., the difference is on the same level as the impact of a different loss function) if the specific data splits are not identical. To address this, we evaluate all methods on identical data splits. In this section, we discuss the used datasets and the data-splitting strategy.
Datasets
We evaluate the methods using a total of 7 datasets: AgeDB [20], AFAD [21], CACD2000 [4], CLAP2016 [1], FG-NET [15], MORPH [23], and UTKFace [28]. We also utilize the IMDB-WIKI dataset [24] for pre-training. However, it is important to note that the labels in the IMDB-WIKI dataset are known to be noisy. To mitigate this problem, Lin et al. [19], and Franc and Čech [11] attempted to clean the labels. To assess the quality of these labels, we trained ResNet-50 models on the dataset using the labels and evaluated the models’ performance on the other datasets. The results are presented in Tab. 1. Both models achieved similar results, so the choice of labels between [19, 11] is arbitrary. Due to a slightly lower overall Mean Absolute Error (MAE), we have decided to use the labels from Franc and Čech [11] for pre-training in the rest of this paper. We further analyze the importance of pre-training in Sec. 4.2.
Data Splits
For the CLAP2016 and CACD2000 datasets, we use the single data split provided by the dataset authors. However, for the remaining datasets, we create 5 subject-exclusive data splits, ensuring that each split has the same age distribution. In this setup, 60% of the dataset is used for training, 20% for model selection, and 20% for evaluating the model performance. Due to its small size, we only use FG-NET for evaluation.
4 Experiments
We conducted experiments to compare various deep-learning methods for age estimation. We also evaluated how different factors, such as data preparation, architecture, and pre-training, affect the performance of the models. In Sec. 4.1 we provide a comparative analysis of methods that were pre-trained on the IMDB-WIKI dataset and then fine-tuned on downstream datasets. The importance of the pre-training is assessed in Sec. 4.2. We also investigate how the choice of the backbone architecture impacts model performance in Sec. 4.3. Lastly, we evaluate different components of the data preparation pipeline in Sec. 4.4. Specifically, we examine the influence of the facial alignment, level of facial coverage, image resolution, and image representation.
4.1 Method Comparison
To fairly compare the performance of the methods, we start their training from the same initialization, specifically from weights pre-trained on ImageNet and then further pre-trained on IMDB-WIKI with cross-entropy. After the pre-training, the last layer of the model is replaced with a layer specific to the desired method. The models are then fine-tuned on the downstream dataset. It is important to note that for the baseline cross-entropy, we also replace the final layer prior to fine-tuning. This ensures that the experimental setup remains identical to that of the other methods. The performance of the models, evaluated using the Mean Absolute Error (MAE), is presented in Tab. 8.
Intra-Dataset Performance
The intra-dataset results achieved with the IMDB-WIKI pre-training are highlighted with a grey background. To determine whether any method is consistently better than others, we employ the Friedman test and the Nemenyi critical difference test as described by Demšar [6]. Using a significance level (p-value) of $\alpha=5\%$, we can conclude that there is no significant difference in performance between any of the methods [21, 12, 13, 9, 22, 17] and the cross-entropy method. In other words, we do not observe any systematic improvement by using any of the methods.
Cross-Dataset Generalization
Cross-dataset results, Tab. 8, were obtained by assessing the performance of models on datasets that were not part of their training. For all of the methods, the cross-dataset performance is consistently and significantly worse than the intra-dataset performance. It is important to note that for computing the cross-dataset performance, the entire evaluation dataset is utilized as the test set. Using the Friedman test and the Nemenyi critical difference test with a significance level (p-value) of $\alpha=5\%$, we can conclude that there is no significant difference in generalization capability between any of the methods [21, 12, 13, 9, 22, 17] and the cross-entropy.
4.2 Importance of Pre-training
We also investigate the impact of pre-training data on the results. We compare the following three scenarios: (i) models initialized with random weights (ii) models initialized with weights pre-trained on ImageNet (iii) models initialized with weights pre-trained on ImageNet and further pre-trained on IMDB-WIKI . The results are presented in Tab. 8, where the initialization scenarios are labeled as follows: (i) Rand. (ii) Imag. (iii) IMDB .
Random Initialization
When starting from random initialization, we noticed that training with the Unimodal loss [17] tends to be unstable. To draw conclusions, we again use the Friedman test and the Nemenyi test with a significance level (p-value)of $\alpha=5\%$, excluding the Unimodal loss [17]. The results indicate that with the random initialization, OR-CNN [21], DLDL [12], and the Mean-Variance loss [22] demonstrate a significant performance improvement over the cross-entropy.
In scenarios with limited data availability, where pre-training is not possible, it is, thus, advisable to utilize one of the aforementioned methods.
Pre-training Data
Our findings reveal that the IMDB-WIKI pre-training method consistently outperforms other approaches in terms of performance. Nonetheless, in certain isolated cases, ImageNet initialization has shown superior results. To draw definitive conclusions, we use the Friedman test and the Nemenyi test, with a significance level (p-value) of $\alpha=5\%$. Our analysis indicates that pre-training with IMDB-WIKI consistently outperforms pre-training solely on ImageNet, and both methods outperform random initialization.
Pre-training Loss Function
It can be argued that the utilization of cross-entropy for the IMDB-WIKI pre-training puts the other methods at a disadvantage, as the extracted features become more suitable for cross-entropy rather than for the alternative methods. In order to investigate this possibility, we have chosen the two most recent methods, namely the Mean-Variance loss proposed by Pan et al. [22] and the Unimodal loss presented by Li et al. [17]. We pre-train the models on IMDB-WIKI using the corresponding loss functions and subsequently finetune them on the downstream datasets. The obtained results are displayed in Tab. 6. We do not find any benefit when using the same method for both pre-training and finetuning. On the contrary, we observe that the Unimodal loss [17] yields better results across all datasets when the model is pre-trained using cross-entropy. We, therefore, conclude that the pre-training with cross-entropy does not invalidate our results.
4.3 Model Architecture
In the age estimation literature, numerous backbone architectures can be found. We aim to evaluate the influence of the architecture choice on the obtained results and present our findings in Tab. 7. We observe that the selection of the model has a more pronounced impact on the performance than the choice of the age estimation method itself.
@@ 4.4 Data Preparation Pipeline
Age estimation models require only a specific region of an image, specifically the person’s face, as input, rather than the entire image. However, the influence of this selection process on the model’s performance is not apriori known. Additionally, facial images used for age estimation can differ in terms of scale and resolution since they originate from various sources and as such need to be resized to a uniform resolution. In this section, we examine the impact of the aforementioned data preparation pipeline on the performance of age estimation models.
@@
Numerous studies lack an explanation of their facial alignment procedure. Others merely mention the utilization of facial landmarks. To assess whether a standardized alignment is needed for a fair comparison of multiple methods, we adopt three distinct alignment procedures and evaluate their effect on model performance. Firstly, we (i) perform no alignment and employ the bounding box proposed by the facial detection model [8] as the simplest approach. Secondly (ii) we utilize the proposed bounding box but rotate it to horizontally align the eyes. Lastly (iii) we use our baseline alignment procedure, which normalizes the rotation, positioning, and scale, and is described in Sec. 3. A visual representation of these facial alignment methods is depicted in Fig. 2. The performance of models trained using the various alignment procedures is presented in Tab. 2. When working with pre-aligned datasets like AFAD, we observe that procedure (iii) does not yield significant improvements compared to the simpler variants (i) or (ii). Similar results are obtained on datasets collected under standardized conditions, such as the MORPH dataset. However, when dealing with in-the-wild datasets like AgeDB and CLAP2016, we find that alignment procedure (iii) leads to noticeable improvements over the simpler methods. Interestingly, on the UTKFace dataset, which also contains in-the-wild images, approach (ii) of solely rotating the proposed bounding boxes achieves superior outcomes compared to approach (iii). In summary, the disparities among the various alignment procedures are not substantial. Hence, it can be argued that any facial alignment technique that effectively normalizes the position, rotation, and scale of the faces would yield comparable results.
@@
We also investigate the impact of facial coverage on the performance of age estimation models. For a visual representation of different levels of facial coverage, refer to Fig. 3. The performance of models trained with the different coverage levels is presented in Tab. 3. Generally, complete facial coverage, which includes the entire head in the model input, yields the best results across the majority of datasets. However, in certain cases such as the AFAD dataset or the MORPH dataset, partial coverage performs better. It is important to note that the AFAD dataset contains preprocessed images that do not capture the entire head. Consequently, using complete facial coverage with this dataset results in the presence of black bars and a decrease in the effective pixel resolution of the face. Therefore, it is expected that increased facial coverage would yield inferior results compared to partial coverage on this dataset. The smallest coverage, limited to the facial region up to the eyes and mouth, consistently performs the worst. It can be concluded that with sufficient pixel resolution, the full facial coverage is superior.
@@
To investigate the influence of input resolution on age estimation, we performed experiments using multiple resolutions on all datasets: specifically, $256\times 256$, $128\times 128$, and $64\times 64$ pixels. The performance of the models trained with different input resolutions is presented in Tab. 4. Our findings indicate that an increase in image resolution consistently results in improved model performance across all datasets. Hence, the best performance was achieved with a resolution of $256\times 256$ pixels. However, it remains uncertain whether further increases in image resolution would yield even more impressive results.
@@
We also examined the input transformation proposed by Lin et al. [18], which involves converting a face image into a tanh-polar representation. This approach has shown great performance in face semantic segmentation. Lin et al. then modified the network for age estimation, reporting impressive results [19]. We explored the potential benefits of applying this transformation for age estimation. However, our findings indicate that the transformation does not improve the results compared to the baseline, as shown in Tab. 5. Therefore, we conclude that the improved age estimation performance observed by Lin et al. [19] does not arise from the use of a different representation, but rather from pre-training on semantic segmentation or their specialized model architecture.
@@
Evaluation
Dataset
Annotations
EM-CNN [11] FP-AGE [19]
AgeDB 6.44 6.30
AFAD 6.86 7.23
CACD2000 5.81 5.90
CLAP2016 6.24 5.53
FG-NET 10.32 6.09
MORPH 4.94 5.30
UTKFace 8.31 6.26
Overall 6.28 6.36
IMDB 4.90 5.15
@@
Table 5: MAE $\downarrow$ of ResNet-50 models with different input transformations. The models were pre-trained on IMDB-WIKI.
@@
Transform
Cross-E. Mean-Var. [22] Unimod. [17]
AgeDB 5.81 5.75 6.17
AFAD 3.14 3.18 3.31
CACD2000 3.96 4.06 4.23
CLAP2016 4.49 4.30 5.25
MORPH 2.81 2.83 2.91
UTKFace 4.38 4.49 4.66
Table 6: Intra-dataset MAE $\downarrow$ of different methods. The models were pre-trained on IMDB-WIKI. Pre-training and finetuning were performed with the same method.

### v3

@@ 5.1 Model Architecture
Multiple different backbone architectures can be found in the age estimation literature. Among these architectures, VGG16 [14, 29, 28, 23, 18, 10] and ResNet-50 [3, 2, 20] stand out as the most common choice. We evaluate the influence of the architecture choice on the performance and extend the comparison to include more recent advancements, EfficientNet-B4 and ViT-B-16. We present our findings in Tab. 5. No backbone emerges as universally best across all datasets. Notably, changes in the backbone have a more substantial impact on performance than changes to the loss function, see Tab. 6. This highlights the importance of a thorough ablation, as changes in the backbone architecture could obscure the impact of the loss function.
@@ 5.2 Data Preparation Pipeline
Age estimation models require only a specific region of an image, specifically the person’s face, as input, rather than the entire image. However, the influence of this selection process on the model’s performance is not apriori known. Should the model be presented with a tight crop of the face or the entire head? Additionally, facial images can differ in terms of scale and resolution since they originate from various sources and as such need to be resized to a uniform resolution. In this section, we examine the impact of the aforementioned data preparation pipeline on the performance of age estimation models. We demonstrate that changes in the preprocessing have a more substantial impact on performance than changes to the loss function.
@@
Numerous studies lack an explanation of their facial alignment procedure. Others merely mention the utilization of facial landmarks. To assess whether a standardized alignment is needed for a fair comparison of multiple methods, we adopt three distinct alignment procedures and evaluate their effect on model performance. Firstly, we {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, ]
perform no alignment and employ the bounding box proposed by the facial detection model [8] as the simplest approach. The bounding box sides are parallel to the axes of the image. Secondly
we utilize the proposed bounding box but rotate it to horizontally align the eyes. Lastly
we use an alignment procedure, which normalizes the rotation, positioning, and scale. For details, refer to the implementation. A visual representation of these facial alignment methods is depicted in Fig. 2. The performance of models trained using the various alignment procedures is presented in Tab. 1. When working with pre-aligned datasets like AFAD, we observe that procedure (iii) does not yield significant improvements compared to the simpler variants (i) or (ii). Similar results are obtained on datasets collected under standardized conditions, such as the MORPH dataset. However, when dealing with in-the-wild datasets like AgeDB and CLAP2016, we find that alignment (iii) leads to noticeable improvements over the simpler methods. Interestingly, on the UTKFace dataset, which also contains in-the-wild images, approach (ii) of solely rotating the proposed bounding boxes achieves the best outcomes. However, the disparities among the various alignment procedures are not substantial. We therefore argue that any facial alignment technique that effectively normalizes the position, rotation, and scale of the faces would yield comparable results.
@@
While facial alignment defines the positioning, orientation, and scale of facial landmarks, the extent to which the face is visible in an image also needs to be specified. We refer to this notion as facial coverage. It measures how much of the face is shown in an image and can range from minimal coverage, where only the eyes and mouth are visible, to complete coverage, where the entire head is visible. Determining the optimal compromise between complete facial coverage and minimal coverage is not immediately clear. Complete facial coverage provides a comprehensive view of the face, allowing age estimation algorithms to consider a broader range of facial cues. On the other hand, partial coverage may help reduce overfitting by eliminating irrelevant facial cues and features with high variance. For a visual demonstration of various facial coverage levels, refer to Fig. 3. The concept of facial coverage has received limited attention in age estimation literature. Consequently, the extent of facial coverage utilized in previous studies can only be inferred from the images presented in those works. For instance, Berg et al. [2] seemingly employ minimal coverage, showing slightly more than just the mouth and eyes. The majority of other works [22, 28, 13, 14, 3, 18] tend to adopt partial coverage, where a significant portion of the face, including the chin and forehead, is visible, but not the entire head and hair. In the works of Pan et al. [23], Rothe et al. [25], and Zhang et al. [29], the entire head is shown.
The performance of models trained with the different coverage levels is presented in Tab. 2. Generally, complete facial coverage, which includes the entire head in the model input, yields the best results across the majority of datasets. However, specifically for AFAD dataset and the MORPH dataset, partial coverage performs better. It is important to note that the AFAD dataset contains preprocessed images that do not capture the entire head. Consequently, using complete facial coverage with this dataset results in the presence of black bars and a decrease in the effective pixel resolution of the face. It is then to be expected that increased facial coverage yields inferior results. The smallest coverage, limited to the facial region up to the eyes and mouth, consistently performs the worst. With sufficient pixel resolution, the full facial coverage performs the best.
@@
To investigate the influence of input resolution on age estimation, we performed experiments using multiple resolutions on all datasets: specifically, $256\times 256$, $128\times 128$, and $64\times 64$ pixels. The results are presented in Tab. 3. Our findings indicate that an increase in image resolution consistently results in improved model performance across all datasets. Hence, the best performance was achieved with a resolution of $256\times 256$ pixels.
In the literature, one can find resolutions ranging from $60\times 60$ to $256\times 256$ pixels, where newer works tend to use larger resolution images. As the resolution increase can directly be observed to improve the results; and the resolutions increased with years; it is difficult to say whether newly proposed methods are better overall, or whether they perform better due to using higher resolution images.
@@
Finally, we examined the input transformation proposed by Lin et al. [19], which involves converting a face image into a tanh-polar representation. This approach has shown large performance improvements in face semantic segmentation. Lin et al. then modified the network for age estimation, reporting impressive results [20]. We explored the potential benefits of applying this transformation for age estimation. However, our findings indicate that the transformation does not improve the results compared to the baseline, as shown in Tab. 4. Therefore, we conclude that the improved age estimation performance observed by Lin et al. [20] does not arise from the use of a different representation, but rather from pre-training on semantic segmentation or their model architecture.
@@
Table 4: MAE $\downarrow$ of ResNet-50 models with different input transformations. The models were pre-trained on IMDB-WIKI [26].
@@
Backbone
ResNet-50 Eff.Net-B4 ViT-B-16 VGG-16
AgeDB
5.81 5.76 9.07 6.02
AFAD
3.14 3.20 4.04 3.22
CACD2000
3.96 4.00 6.22 3.92
CLAP2016
4.49 4.06 8.55 4.65
MORPH
2.81 2.87 4.35 2.88
UTKFace
4.38 4.23 6.88 4.64
5.3 FaRL Backbone
We observed that adjustments to the decision layer and loss function have minimal impact on the final model performance. Conversely, large performance disparities arise when modifying other components of the prediction pipeline. Notably, the pretraining data appear to be the most influential factor. Based on this insight, we opt against creating a specialized loss function to enhance the age estimation system. Instead, we leverage the FaRL backbone by Zheng et al. [31], utilizing a ViT-B-16 [9] model. The FaRL model is trained through a combination of (i) contrastive loss on image-text pairs and (ii) prediction of masked image patches. Training takes place on an extensive collection of facial images (50 million) from the image-text pair LAION dataset [27]. We retain the feature representation extracted by FaRL without altering the model’s weights. Our decision to use FaRL is driven solely by the extensive amount of pre-training data it incorporates, rather than specific characteristics of the backbone. Different image encoders could be trained in the same manner. However, due to the costs associated with training such models, we have chosen to use the available FaRL ViT-B-16 backbone. We employ a simple multilayer perceptron (MLP) over the FaRL-extracted features, consisting of 2 layers with 512 neurons each, followed by ReLU activation. Cross-entropy serves as the chosen loss function. For each downstream dataset, we pretrain the MLP on IMDB-WIKI or initialize it to random weights. We choose the preferred option based on validation loss on the downstream dataset. As previously, we replace the final layer before fine-tuning on downstream datasets.
This straightforward modification outperformed all other models on AgeDB, CLAP2016, and UTKFace datasets. It also achieved superior results on AFAD, matched the performance of other models on CACD2000, but demonstrated worse performance on MORPH. Applying the FN test revealed statistically significant improvements of this model over others in both intra-dataset and cross-dataset evaluations, see Tab. 6. We attribute the poor performance of FaRL on MORPH to the fact that the distributions of images in LAION [27] and MORPH [24] are vastly different. As we do not finetune the feature representation of FaRL [31], it is possible that the representation learned on LAION is superior on the other datasets but deficient on MORPH.
We do not claim this model to be the ultimate solution, but the results achieved with the FaRL backbone along with our public implementation offer a robust and straightforward baseline for a comparison with future methods.